Tuesday, November 18: DAY Shift
From Hall A Wiki
Go back to Daily Run Plan Daily Run Plans
- General considerations
- Vincent Sulkosky is the current Run Coordinator. Please contact me directly via RC phone 757-876-1787
- Follow the generic Instructions for shift takers
- Shift leaders: Please make sure to fill out the beam time accounting form and write a "Shift Summary" log entry at the end of your shift regardless if there is beam or not.
- In general, technical staff should not be contacted by the shift crew - please call the RC first.
- If the run plan calls for some significant period of beam at 10 uA or greater please call Mahmoud Ahmad, who would like to perform a parasitic beam diagnostic test (no impact to us). His contact information will be written on the white board in the Counting House.
- Instructions for this shift
- The target operator is required to supervise the cryo-target.
- Kinematics, DVCS Kin1: Ebeam = 7.3 GeV, PLHRS=1.1981 GeV, Theta_HRS=22.83, Theta_calo=12.35
- Make sure to use the DVCS configuration for all runs.
- Due to continued beam trips at 1C12 and the required work needed to fix the beam issues, Hall A will not receive CW the rest of today. No beam is expected on DAY shift.
- This morning Hall A went to controlled access, primarily to address an issue with the left HRS dipole controls. One this work is completed, Hall A will go back to beam permit.